3 workers take 10 days to make a boat. How long would 5 workers take.
Found 3 solutions by stanbon, jorel1380, lwsshak3:
Answer by stanbon(75887)

3 workers take 10 days to make a boat. How long would 5 workers take.
---
# of workers an # of days are inversely related.
----
d = k/w
--
Solve for "k" using "3 workers take 10 days to make a boat".
10 = k/3
k = 30
-----
Equation:
d = 30/w
---
How long would 5 workers take.
d = 30/5
d = 6 days (# of days for 5 workers to do the job)
----
Cheers,
Stan H.

Answer by jorel1380(3719)

1 worker=1/3x1/10 boats per day=1/30 boats per day
5 workers =5/30 boats per day=1/6 boat per day= 6 days per one boat..

Answer by lwsshak3(11628)

3 workers take 10 days to make a boat. How long would 5 workers take.
**
Assuming all workers have the same individual work rates=1/x
The work rate of 3 workers working together=1/10
1/x+1/x+1/x=1/10
3/x=1/10
x=30
individual work rate =1/30, that is, each worker would take 30 days to complete the job working alone.
With 5 workers working together:
1/30+1/30+1/30+1/30+1/30=1/x
5/30=1/x
5x=30
x=6
ans:
It would 5 workers 6 days to make a boat.